OpenStudy (anonymous):

If the sequence is geometric, give its common ratio. 1, 5, 15, 22 ...

OpenStudy (campbell_st):

If a sequence is geometric it has a common ratio to find the common ratio compare the terms \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}} ....... \] so you have \[\frac{5}{1} = \frac{15}{5} = \frac{22}{15}\] the ratios are not the same so the sequence isn't geometric